OpenStudy (anonymous):

Kate is three years less thantwice as old as her brother,Brett. Next year Brett's age will be three-fourths Kate's age. How old is each now?

OpenStudy (anonymous):

Let K=Kate and B=Brett. K-3=2B K-3+1=2B+1 K-2 = 2B+1 K-3 = 2B (K-3)/2 = B or K= 2B+3 We also know that K=(3/4 * B) + 1 K = 2B + 3 is same as (3/4 * B) + 1 Therefore K - 1 = 2B+2 = (3/4 * B) ...

OpenStudy (anonymous):

Substitution with (K-3)/2 = B or K= 2B+3

OpenStudy (anonymous):

I on't understand

OpenStudy (anonymous):

Subbed (3/4 * B) + 1 = K's age, with (3/4*(K-3)/2) + 1 = K 3k-9/8 = K So, 8k-3k = -9 + 8 3K = -1 Since we know that next year brett age is 3K = -1, then now K=1/3 of Brett age